OpenStudy (steve816):

Help me with some algebra 2 ! I'm suffering from summer learning loss :(

OpenStudy (steve816):

Given \[f(x)=\frac{ x^2+7x+12 }{ x^2-16 }\] Find the folloving: x intercept(s) y intercept equation(s) of the vertical asymptote Domain Equation of the horizontal/oblique asymptote location of hole(s) in the graph end behavior

OpenStudy (steve816):

I don't need you guys to find the answers I need you guys to refresh my memory and explain how to find each one! Help would be greatly appreciated.

OpenStudy (campbell_st):

so to find the x- intercepts solve \[x^2 + 12x + 12 = 0\] but check its not an asymptote... to find the y-intercepts let x = 0 and solve for y vertical asymptotes are the place where the denominator is zero, so solve \[x^2 - 16 = 0\] Domain, what values of x, can't you have in the denominator..? the curve has a horizontal asymptote since the degree of the numerator and denominator are both 2. so simplify \[\frac{x^2}{x^2} = ?\] and if you get stuck just graph it using https://www.desmos.com/calculator

OpenStudy (steve816):

Man thanks so much. You know everything xD